A backyard has dimensions 1 1/3 yards by 3 4/5 yards. What is the area of the back yard in square yards (yd2)?
Answer by edjones(7311)

L*w=A
4/3 * 19/4 = 76/12 = 19/3 = 6 1/3 yd^2